# pQT_AMTs.bib

### v1

@@ 5.1 Lüders measurements: quantum theory
Another argument to extend the single-system Lüders rule to composite systems uses the fact that $\omega^{\mathsf{L}}_{A,x}(\cdot)\equiv\omega^{\mathsf{L}}_{A}(P^{A}_{x},\cdot):\bar{\mathcal{S}}(\mathcal{H}_{A})\to\bar{\mathcal{S}}(\mathcal{H}_{A})$, which denotes the Lüders rule for system $\mathcal{H}_{A}$ conditioned on the measurement outcome $P_{x}^{A}$, is convex-linear over $\bar{\mathcal{S}}(\mathcal{H}_{A})$. Therefore, $\omega^{\mathsf{L}}_{A,x}$ has a unique linear extension to the set of bounded operators $\mathcal{L}(\mathcal{H}_{A})$ (cf. [heinosaari_mathematical_2011]). The Lüders rule (15) for the composite system $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$, conditioned on the local measurement outcome $P_{x}^{A}$, is then obtained by setting
@@ 5.3 Non-collapsing measurements: passive quantum theory
We will now consider measurements that, instead of updating the quantum state, leave it unchanged. [^10] The possibility of non-collapsing measurements was listed by von Neumann [vonneumann_mathematische_1932] as early as 1932 as one of three possible reactions of a physical system to a measurement. Outcomes still occur probabilistically as dictated by the Born rule, yet, in analogy with classical theory, these measurements do not disturb the pre-measurement state. This situation corresponds to a “passive” update rule $\omega^{\mathsf{P}}_{A}$ for non-composite systems,
@@
Non-collapsing measurements, originally introduced in [aaronson_space_2016], cannot be implemented in quantum theory [busch_is_1997]. Indeed, $\omega^{\mathsf{P}}_{A}$ is not convex-linear over $\bar{\mathcal{S}}(\mathcal{H}_{A})$ which implies that, after a passive measurement, one may be able to distinguish between different preparations of the same mixed state. For example, let $\mathcal{H}_{A}=\mathbb{C}^{2}$ and consider the Gemenge
@@
where |±$|\pm\rangle=(|0\rangle\pm|1\rangle)/\sqrt{2}$. Both preparations correspond to the maximally mixed state, $\rho_{\mathcal{G}}=\rho_{\mathcal{G}^{\prime}}=\mathbb{I}_{2}/2$. According to the Gemenge-update rule (14) in Sec. 4.3, a passive measurement with outcome $P_{x}^{A}=|0\rangle\langle 0|$ will produce two different sub-normalised density operators,
@@
While complete positivity can be defined for both linear and nonlinear transformations [ando_non-linear_1986], it has been described as “physically unfitting” for nonlinear dynamics [czachor_complete_1998]. In fact, although $\omega^{\mathsf{P}}_{A}$ (or, more precisely, extensions of $\omega^{\mathsf{P}}_{A}$ to $\mathcal{L}(\mathcal{H}_{A})$) is not completely positive, the update rule can still be consistently extended to composite systems. Invoking compositional compatibility (Def. 3), we set
@@
The map $\omega^{\mathsf{P}}$ satisfies all the conditions in Def. 1 and defines a GUT in which measurements do not modify the state of either the measured system or of the larger composite system to which it belongs. We called this theory passive quantum theory (pQT). An overview of pQT and its key properties has been provided in [fiorentino_quantum_2023].
The absence of any measurement-induced state update allows, in principle, for tomographic reconstruction of the state of an individual system. Thus, the cloning of arbitrary states becomes possible and makes computation in pQT more efficient (in some sense) than in standard quantum theory [busch_is_1997, kent_nonlinearity_2005]. The passive update rule of pQT also satisfies ideality but is not locally tomographic and does not allow for violations of Bell’s inequalities (cf. Table 1).
@@ 5.6 Invalid state-update rules
The projection postulate introduced by von Neumann [vonneumann_mathematische_1932] is a notable example of state-update that does not define a valid update rule. For non-degenerate measurements, von Neumann’s and Lüders’ postulates coincide. However, according to von Neumann, outcome degeneracy of quantum measurements—represented by projectors of rank greater than one—arises from classical post-processing of non-degenerate measurements. That is, one does not directly implement a measurement of a degenerate observable $M$; instead, a measurement of a refinement of $M$, i.e. a non-degenerate observable $M^{\prime}$ commuting with $M$, is carried out, and the outcomes are coarse-grained. Given a degenerate observable, there exist, however, infinitely many refinements. Consequently, the post-measurement state will not only depend solely on the pre-measurement state $\rho_{A}$ (which we can assume to be pure, without loss of generality) and the observed degenerate outcome $P^{A}_{x}$, but also on the specific refinement chosen. More precisely, letting $\delta=\text{Tr}{(P^{A}_{x})}$ denote the “cardinality” of the degeneracy of outcome $P^{A}_{x}$, the updated state depends on the subset $\mathcal{B}_{x}=\{P^{A}_{x_{i}}\}_{i=1}^{\delta}\subseteq M^{\prime}$ spanning the degenerate subspace $\mathcal{H}_{x}=P^{A}_{x}\mathcal{H}$. The state-update for non-composite systems thus obeys
@@
Finally, we present a simple example of an update rule that violates local covariance A4. Assume that any measurement on a system causes the pre-measurement state to undergo a fixed, observer-independent and non-trivial [^14] If $U=\mathbb{I}$, the passive measurements of Sec. 5.3 result consistent with A4. unitary transformation $U\in\mathcal{U}(\mathcal{H})$,
@@ 6 How to recover the projection postulate
Lüders presents two main objections to von Neumann’s projection postulate [luders_uber_1950]. Firstly, he argues that the post-measurement state should depend only on the outcome and the initial state. This requirement is equivalent to context-independence A2. Secondly, he suggests that
“the measurement of a highly degenerate quantity permits only relatively weak assertions regarding the considered ensemble. For that reason, the resulting change in state should likewise be small” ([luders_concerning_2006], p. 665).
@@
In some form or another, the assumption of coherence has appeared already in the literature, both in the context of quantum theory [bell_moral_1966, herbut_compatibility_2004, khrennikov_neumann_2009] and of generalised probabilistic theories [kleinmann_sequences_2014]. For measurements on non-composite quantum systems, coherence entails both deterministic repeatability [kleinmann_sequences_2014] and ideality.
@@
Indeed, coherence alone is sufficient to derive the single-system Lüders rule $\omega^{\mathsf{L}}_{A}$, i.e. the original projection postulate ($\mathsf{L}$). This argument was first presented in 1966 [bell_moral_1966] and similar versions of it have since been rediscovered several times; see Appendix A for a summary. Below, we present a simple proof of this result.
@@
Additional assumptions are necessary to fully recover the Lüders rule $\omega^{\mathsf{L}}_{AB}$ for composite systems. One such assumption is composition compatibility (Def. 3)—see Theorem 2 below. This final step in deriving the projection postulate is often overlooked [bell_moral_1966, herbut_compatibility_2004, martinez_search_1990], as the extension of $\omega^{\mathsf{L}}_{A}$ to local measurements on composite systems is typically assumed implicitly, despite not being stated in or implied by the standard postulates.
@@ 7.1 Summary
Any valid state-update rule must satisfy six core requirements which include the uniqueness of post-measurement states, no-signalling, and consistency for measurements carried out by observers with access to subsystems only. We have shown that quantum theory and passive quantum theory [fiorentino_quantum_2023] satisfy the requirements. Three new examples of consistent generalised state-update rules have been defined, each one leading to a toy theory structurally different from quantum theory.
@@
Having established a variety of acceptable update rules, it is natural to look for properties that single out the Lüders rule. We have shown that the property of coherence uniquely implies the Lüders rule for non-composite systems. This result agrees with earlier ones based on similar concepts describing a form of information-disturbance trade-off. However, our systematic approach to GURs highlights the fact that known derivations of the Lüders rule are incomplete since they typically do not take composite systems into account. We find that coherence on its own is insufficient to single out the projection postulate when considering composite systems; equally, composition compatibility alone is not strong enough to do so. However, combining it with coherence implies the projective Lüders rule for composite systems, too (Theorem 2).
@@ 7.2 Discussion and outlook
Non-linear update rules may enable observers to distinguish between proper and improper mixtures. In the corresponding foil theories, density operators will not provide a complete description of the hypothetical physical systems. Importantly, complete positivity is not strong enough to ensure that such measurement-induced transformations assign post-measurement states to composite systems consistently. In the words of [czachor_complete_1998], complete positivity is “physically unfitting” for this purpose.
Reconstructions of quantum theory that engage—either directly or indirectly—with post-measurement states usually assume measurement-induced state transformations to be linear in the pre-measurement state [barrett_information_2007, kleinmann_sequences_2014, masanes_measurement_2019]. This assumption, often left implicit [stacey_masanes-galley-muller_2022, kent_measurements_2023], excludes a priori certain types of measurement behaviours—such as the trivial non-collapsing rule of “passive quantum theory”—that are, in fact, operationally well-defined and consistent with fundamental principles such as no-signalling (see Sec. 4.2). Preliminary results based on the framework of GURs (as well as earlier works, e.g. [flatt_gleason-busch_2017]) suggest that the linearity of update rules could be derived from other, more natural operational assumptions.
@@ Acknowledgements
The authors are grateful for support through the Leverhulme grant RPG-2024-201.
Appendix A Summary of derivations of the Lüders rule
@@
In 1966, Bell and Nauenberg showed that the quantum mechanical state-update can be obtained by assuming that the outcome probability distribution of a common refinement for commuting observables $M_{1}$ and $M_{2}$ equals the joint probability distribution obtained by measuring $M_{1}$ and $M_{2}$ sequentially, in any order [bell_moral_1966]. This assumption is mathematically equivalent to that of coherence (Def. 4) employed in the proof of Theorem 1 in Sec. 6. To the best of our knowledge, the only reference to this early operational derivation of the Lüders rule in non-composite systems is by Herbut [herbut_compatibility_2004]. Over the last 40 years, the argument by Bell and Nauenberg has been rediscovered at least three times, using the concept of coherence.
@@
In 1983, Cassinelli and Zanghì formulated quantum theory in terms of a generalised probability space and showed how to recover the single-system Lüders rule in this context. To do so, they describe the state-update following measurements in terms of a generalised conditional probability defined on the generalised probability space—i.e. a probability measure on $\mathcal{P}(\mathcal{H})$ compatible with Born’s rule and satisfying “coherence” [cassinelli_conditional_1983].
@@
In 2009, Khrennikov showed that the Lüders rule for single systems follows from von Neumann’s projection postulate for non-degenerate observables if “coherence” is assumed [khrennikov_neumann_2009]. Claiming that “this important observation remained unnoticed for the past 70 years”, the author appears to be unaware of Bell and Nauenberg’s work.
@@
In 2014, Kleinmann developed a formalism for sequential measurements in a broad class of generalised probabilistic theories (GPTs), introducing a generalisation of the Lüders rule based on coherence [kleinmann_sequences_2014]. A simple derivation of the non-composite Lüders rule based on coherence is also given, without reference to earlier work. When applied to quantum theory, i.e. in a Hilbert-space setting for states, Kleinmann’s approach reproduces convex-linear update rules for single systems. This restriction is due to assuming preparation indistinguishability throughout (see Sec. 5.1), thereby excluding operationally valid alternatives such as the passive update rule from the outset.
@@
Herbut actually presents two distinct derivations applying to measurements on non-composite systems in a paper from 1969 [herbut_derivation_1969]. In an information-theoretic approach, he postulates (i) deterministic repeatability (see Sec. 2.1) and that (ii) the post-measurement state minimises its distance from the pre-measurement state. The Lüders rule is then shown to map the pre-measurement state to the least distinguishable state compatible with outcome repeatability. This result remained largely unnoticed, while anticipating later arguments deriving the state-update from minimisation principles, including those by Marchand [marchand_statistical_1977], Dieks [dieks_distance_1983], and Hadjisavvas [hadjisavvas_distance_1981]. More recently, the single-system Lüders rule was obtained from minimising non-symmetric quantum information distances [kostecki_luders_2014, hellmann_quantum_2016], thereby establishing a mathematical connection to the Bayes-Laplace rule. Ozawa [ozawa_quantum_1984, ozawa_quantum_1998] derives the update rule from Bayes’ rule by assuming the standard joint probability formula for local quantum measurements.
@@
Herbut’s second derivation in [herbut_derivation_1969] takes an operational approach. In addition to deterministic repeatability, it is assumed that performing an unconditional measurement of observable $M$ does not affect the expectation values of any observables commuting with $M$. The argument focuses on measurements on single systems. However, Herbut also briefly considers composite systems, discussing in particular how local Lüders measurements modify joint states without enabling signalling. When doing so, he implicitly adopts the standard extension to composite systems (see Sec. 5.1). What is more, Herbut takes convex-linearity over states to be a fundamental requirement for state updates, without justifying this assumption.
@@
Herbut revisits the question in a 1974 paper [herbut_minimal_1974]. Here, the Lüders rule for single systems is obtained by combining the assumption of deterministic repeatability with the “preservation of sharp values”. An observable $M$ has a sharp value $x$ if $\text{Tr}(P_{x}^{A}\rho)=1$ holds for some $P_{x}\in M$. This value is preserved if, given an initial state $\rho$ with a sharp $M$-value $x$, a measurement of any observable commuting with $M$ leaves unchanged the value $x$, regardless of the observed outcome. When proving this result for improper mixtures, Herbut implicitly adopts “composition compatibility” to describe local measurements.
@@
Martinez [martinez_search_1990] offers an early account of known derivations of the Lüders rule including Herbut’s work but not Bell and Nauenberg’s contribution. The paper also demonstrates that the converse of Lüders theorem [luders_uber_1950, busch_quantum_1996] implies the quantum mechanical collapse for non-composite systems: if, for any pair of commuting observables $M_{1}$ and $M_{2}$ and any initial state, we assume that the sequence of measurements $M_{1}\prec M_{2}\prec M_{1}$ always produces identical outcomes for $M_{1}$ regardless of the outcome of the $M_{2}$ measurement, then the post-measurement state must necessarily be given by the Lüders rule. Martinez further examines a quantum-logical derivation employing the “Sasaki hook” presented in [friedman_quantum_1978]. The extension of the update rule to local measurements on subsystems is not considered.
@@
The proof that establishes the Lüders instrument as the unique ideal quantum instrument for sharp discrete measurements dates back to Davies [davies_quantum_1976], as recognised in [busch_quantum_1996, busch_repeatable_1995]. The notion of a quantum instrument incorporates two key assumptions: convex-linearity of the single-system state-update rule, and the standard linear extension to composite systems enabled by the complete positivity of $\omega_{A}$ (see Eq. (18) in Sec. 5.1). This argument for isolating the Lüders collapse for single systems has been rediscovered occasionally (see  [marinkovic_note_1983], for example).
@@
In 2019, Masanes et al. advance the argument that both the Born rule and the post-measurement state rule can be derived from the other postulates of quantum theory when supplemented by suitable operational constraints [masanes_measurement_2019]. In doing so, they consider alternative rules for computing outcome probabilities, explicitly discussing both composite systems and local measurements. Having established the uniqueness of the Born rule within their framework, they assert that quantum instruments constitute the only consistent description of state updates. This conclusion, along with their broader framework, has attracted some criticism [stacey_masanes-galley-muller_2022, galley_reply_2022, kent_measurements_2023, masanes_response_2023]. The argument for quantum instruments depends on an assumption regarding sequential measurements that, as already shown in [flatt_foundations_2019], is mathematically equivalent to requiring convex-linearity of state updates.
@@
Conceptually, Masanes et al.’s approach mirrors Kleinmann’s framework [kleinmann_sequences_2014] in that it is limited a priori to update rules satisfying preparation indistinguishability. This restriction explains why they adopt the completely positive extension (18) for local measurements on composite systems: any other extension would necessarily violate their assumption of preparation indistinguishability.
@@
In retrospect, the Lüders rule is seen to have been derived on the basis of operational, information-theoretic, quantum-logical, and epistemically motivated assumptions. Most of the derivations are limited to single-system update rules, and their extension to local measurements on subsystems is either ignored or tacitly assumed. The extension of state-update rules is, however, not trivial, since the single-system rule alone does not determine it. Additional structural assumptions such as preparation indistinguishability or composition compatibility (cf. Theorem 2) are required to exclude alternative update rules in the multi-partite case. None of these properties are contained in the postulates ($\mathsf{S}$) to ($\mathsf{B}$) of quantum theory.
[…]
[heading=bibintoc]

### v2

@@ 5.1 Lüders measurements: quantum theory
Another argument to extend the single-system Lüders rule to composite systems uses the fact that $\omega^{\mathsf{L}}_{A,x}(\cdot)\equiv\omega^{\mathsf{L}}_{A}(P^{A}_{x},\cdot):\bar{\mathcal{S}}(\mathcal{H}_{A})\to\bar{\mathcal{S}}(\mathcal{H}_{A})$, which denotes the Lüders rule for system $\mathcal{H}_{A}$ conditioned on the measurement outcome $P_{x}^{A}$, is convex-linear over $\bar{\mathcal{S}}(\mathcal{H}_{A})$. Therefore, $\omega^{\mathsf{L}}_{A,x}$ has a unique linear extension to the set of bounded operators $\mathcal{L}(\mathcal{H}_{A})$ (cf. [undefap]). The Lüders rule (15) for the composite system $\mathcal{H}_{A}\otimes\mathcal{H}_{B}$, conditioned on the local measurement outcome $P_{x}^{A}$, is then obtained by setting
@@ 5.3 Non-collapsing measurements: passive quantum theory
We will now consider measurements that, instead of updating the quantum state, leave it unchanged. [^13] The possibility of non-collapsing measurements was listed by von Neumann [undefaaz] as early as 1932 as one of three possible reactions of a physical system to a measurement. Outcomes still occur probabilistically as dictated by the Born rule, yet, in analogy with classical theory, these measurements do not disturb the pre-measurement state. This situation corresponds to a “passive” update rule $\omega^{\mathsf{P}}_{A}$ for non-composite systems,
@@
Non-collapsing measurements, originally introduced in [undef], cannot be implemented in quantum theory [undefp]. Indeed, $\omega^{\mathsf{P}}_{A}$ is not convex-linear over $\bar{\mathcal{S}}(\mathcal{H}_{A})$ which implies that, after a passive measurement, one may be able to distinguish between different preparations of the same mixed state. For example, let $\mathcal{H}_{A}=\mathbb{C}^{2}$ and consider the Gemenge
@@
where $|\pm\rangle=(|0\rangle\pm|1\rangle)/\sqrt{2}$. Both preparations correspond to the maximally mixed state, $\rho_{\mathcal{G}}=\rho_{\mathcal{G}^{\prime}}=\mathbb{I}_{2}/2$. According to the Gemenge-update rule (14) in Sec. 4.3, a passive measurement with outcome $P_{x}^{A}=|0\rangle\langle 0|$ will produce two different sub-normalised density operators,
@@
While complete positivity can be defined for both linear and nonlinear transformations [undefb], it has been described as “physically unfitting” for nonlinear dynamics [undefy]. In fact, although $\omega^{\mathsf{P}}_{A}$ (or, more precisely, extensions of $\omega^{\mathsf{P}}_{A}$ to $\mathcal{L}(\mathcal{H}_{A})$) is not completely positive, the update rule can still be consistently extended to composite systems. Invoking compositional compatibility (Def. 3), we set
@@
The map $\omega^{\mathsf{P}}$ satisfies all the conditions in Def. 1 and defines a GUT in which measurements do not modify the state of either the measured system or of the larger composite system to which it belongs. We called this theory passive quantum theory (pQT). An overview of pQT and its key properties has been provided in [undefag].
The absence of any measurement-induced state update allows, in principle, for tomographic reconstruction of the state of an individual system. Thus, the cloning of arbitrary states becomes possible and makes computation in pQT more efficient (in some sense) than in standard quantum theory [undefp, undefay]. The passive update rule of pQT also satisfies ideality but is not locally tomographic and does not allow for violations of Bell’s inequalities (cf. Table 1).
@@ 5.6 Invalid state-update rules
The projection postulate introduced by von Neumann [undefaaz] is a notable example of state-update that does not define a valid update rule. For non-degenerate measurements, von Neumann’s and Lüders’ postulates coincide. However, according to von Neumann, outcome degeneracy of quantum measurements—represented by projectors of rank greater than one—arises from classical post-processing of non-degenerate measurements. That is, one does not directly implement a measurement of a degenerate observable $M$; instead, a measurement of a refinement of $M$, i.e. a non-degenerate observable $M^{\prime}$ commuting with $M$, is carried out, and the outcomes are coarse-grained. Given a degenerate observable, there exist, however, infinitely many refinements. Consequently, the post-measurement state will not only depend solely on the pre-measurement state $\rho_{A}$ (which we can assume to be pure, without loss of generality) and the observed degenerate outcome $P^{A}_{x}$, but also on the specific refinement chosen. More precisely, letting $\delta=\text{Tr}{(P^{A}_{x})}$ denote the “cardinality” of the degeneracy of outcome $P^{A}_{x}$, the updated state depends on the subset $\mathcal{B}_{x}=\{P^{A}_{x_{i}}\}_{i=1}^{\delta}\subseteq M^{\prime}$ spanning the degenerate subspace $\mathcal{H}_{x}=P^{A}_{x}\mathcal{H}$. The state-update for non-composite systems thus obeys
@@
Finally, we present a simple example of an update rule that violates local covariance A3. Assume that any measurement on a system causes the pre-measurement state to undergo a fixed, observer-independent and non-trivial [^17] If $U=\mathbb{I}$, we recover the passive measurements of Sec. 5.3, which are consistent with A3. unitary transformation $U\in\mathcal{U}(\mathcal{H})$,
@@ 6 How to recover the projection postulate
Lüders presents two main objections to von Neumann’s projection postulate [undefaaf]. Firstly, he argues that the post-measurement state should depend only on the outcome and the initial state. This requirement is equivalent to context-independence A2. Secondly, he suggests that
“the measurement of a highly degenerate quantity permits only relatively weak assertions regarding the considered ensemble. For that reason, the resulting change in state should likewise be small” ([undefaae], p. 665).
@@
In some form or another, the assumption of coherence has appeared already in the literature, both in the context of quantum theory [undefi, undefat, undefaaa] and of generalised probabilistic theories [undefaab]. For measurements on non-composite quantum systems, coherence entails both deterministic repeatability [undefaab] and ideality.
@@
Indeed, coherence alone is sufficient to derive the single-system Lüders rule $\omega^{\mathsf{L}}_{A}$, i.e. the original projection postulate ($\mathsf{L}$). This argument was first presented in 1966 [undefi] and similar versions of it have since been rediscovered several times; see Appendix A for a summary. Below, we present a simple proof of this result.
@@
Additional assumptions are necessary to fully recover the Lüders rule $\omega^{\mathsf{L}}_{AB}$ for composite systems. One such assumption is composition compatibility (Def. 3)—see Theorem 2 below. This final step in deriving the projection postulate is often overlooked [undefi, undefat, undefaai], as the extension of $\omega^{\mathsf{L}}_{A}$ to local measurements on composite systems is typically assumed implicitly, despite not being stated in or implied by the standard postulates.
@@ 7.1 Summary
Any valid state-update rule must satisfy six core requirements which include the uniqueness of post-measurement states, no-signalling, and consistency for measurements carried out by observers with access to subsystems only. We have shown that quantum theory and passive quantum theory [undefag] satisfy the requirements. Three new examples of consistent generalised state-update rules have been defined, each one leading to a toy theory structurally different from quantum theory.
@@
Having established a variety of acceptable update rules, it is natural to look for properties that single out the Lüders rule. We have shown that the property of coherence uniquely implies the Lüders rule for non-composite systems. This result agrees with earlier ones based on similar concepts describing a form of information-disturbance trade-off. However, our systematic approach to GURs highlights the fact that many derivations of the Lüders rule may be considered incomplete by not explicitly addressing composite systems. We find that coherence on its own is insufficient to single out the projection postulate when considering composite systems; equally, composition compatibility alone is not strong enough to do so. However, combining it with coherence implies the projective Lüders rule for composite systems, too (Theorem 2).
@@ 7.2 Discussion and outlook
Non-linear update rules may enable observers to distinguish between proper and improper mixtures. In the corresponding foil theories, density operators will not provide a complete description of the hypothetical physical systems. Importantly, complete positivity is not strong enough to ensure that such measurement-induced transformations assign post-measurement states to composite systems consistently. In the words of [undefy], complete positivity is “physically unfitting” for this purpose.
Reconstructions of quantum theory that engage—either directly or indirectly—with post-measurement states usually assume measurement-induced state transformations to be linear in the pre-measurement state [undefe, undefaab, undefaak]. This assumption, often left implicit [undefaaw, undefax], a priori excludes measurement behaviours such as the trivial non-collapsing rule of “passive quantum theory” that are, in fact, operationally well-defined and consistent with fundamental principles such as no-signalling (see Sec. 4.2). Preliminary results based on the framework of GURs and earlier work (e.g. [undefaan, undefah]) suggest that the linearity of update rules could be derived from natural operational assumptions, mirroring efforts to establish the linearity of the quantum time evolution [undefam, undefaaab].
@@ Acknowledgements
The authors are grateful for support through grant RPG-2024-201 by the Leverhulme Trust.
Appendix A Derivations of the Lüders rule – A brief survey
@@
In 1966, Bell and Nauenberg showed that the quantum mechanical state-update can be obtained by assuming that the outcome probability distribution of a common refinement for commuting observables $M_{1}$ and $M_{2}$ equals the joint probability distribution obtained by measuring $M_{1}$ and $M_{2}$ sequentially, in any order [undefi]. This assumption is mathematically equivalent to that of coherence (Def. 4) employed in the proof of Theorem 1 in Sec. 6. To the best of our knowledge, the only reference to this early operational derivation of the Lüders rule in non-composite systems is by Herbut [undefat]. Over the last 40 years, the argument by Bell and Nauenberg has been rediscovered at least three times, using the concept of coherence.
@@
In 1983, Cassinelli and Zanghì formulated quantum theory in terms of a generalised probability space and showed how to recover the single-system Lüders rule in this context. To do so, they describe the state-update following measurements in terms of a generalised conditional probability defined on the generalised probability space—i.e. a probability measure on $\mathcal{P}(\mathcal{H})$ compatible with Born’s rule and satisfying “coherence” [undeft].
@@
In 2009, Khrennikov showed that the Lüders rule for single systems follows from von Neumann’s projection postulate for non-degenerate observables if “coherence” is assumed [undefaaa]. Claiming that “this important observation remained unnoticed for the past 70 years”, the author appears to be unaware of Bell and Nauenberg’s work.
@@
In 2014, Kleinmann developed a formalism for sequential measurements in a broad class of generalised probabilistic theories (GPTs), introducing a generalisation of the Lüders rule based on coherence [undefaab]. A simple derivation of the non-composite Lüders rule based on coherence is also given, without reference to earlier work. When applied to quantum theory, i.e. in a Hilbert-space setting for states, Kleinmann’s approach reproduces convex-linear update rules for single systems. This restriction is due to assuming preparation indistinguishability throughout (see Sec. 5.1), thereby excluding operationally valid alternatives such as the passive update rule from the outset.
@@
Herbut actually presents two distinct derivations applying to measurements on non-composite systems in a paper from 1969 [undefas]. In an information-theoretic approach, he postulates (i) deterministic repeatability (see Sec. 2.1) and that (ii) the post-measurement state minimises its distance from the pre-measurement state. The Lüders rule is then shown to map the pre-measurement state to the least distinguishable state compatible with outcome repeatability. This result remained largely unnoticed, while anticipating later arguments deriving the state-update from minimisation principles, including those by Marchand [undefaag], Dieks [undefab], and Hadjisavvas [undefan]. More recently, the single-system Lüders rule was obtained from minimising non-symmetric quantum information distances [undefaad, undefaq], thereby establishing a mathematical connection to the Bayes-Laplace rule. Ozawa [undefaan, undefaao] derives the quantum instrument formalism describing state-updates from Bayes’ rule by assuming the standard joint probability formula for local quantum measurements. In a sense, the author adopts a stronger interpretation of postulate $(\mathsf{B})$ (cf. Sec. 3.1) than the one used explicitly by us, and implicitly by other authors [undefay, undef, undefaz, undefax] where $(\mathsf{L})$ is modified. The stronger interpretation of Born’s rule is assumed to govern individual measurement outcomes as well as correlations between local measurements carried out by different experimenters with access to parts of a composite system only. Essentially, it subsumes the property of local tomography (cf. Sec. 5.1).
@@
Herbut’s second derivation in [undefas] takes an operational approach. In addition to deterministic repeatability, it is assumed that performing an unconditional measurement of observable $M$ does not affect the expectation values of any observables commuting with $M$. The argument focusses on measurements on single systems. However, Herbut also briefly considers composite systems, discussing in particular how local Lüders measurements modify joint states without enabling signalling. When doing so, he implicitly adopts the standard extension to composite systems (see Sec. 5.1). What is more, Herbut takes convex-linearity over states to be a fundamental requirement for state updates, without justifying this assumption.
@@
Herbut revisits the question in a 1974 paper [undefau]. Here, the Lüders rule for single systems is obtained by combining the assumption of deterministic repeatability with the “preservation of sharp values”. An observable $M$ has a sharp value $x$ if $\text{Tr}(P_{x}^{A}\rho)=1$ holds for some $P_{x}\in M$. This value is preserved if, given an initial state $\rho$ with a sharp $M$-value $x$, a measurement of any observable commuting with $M$ leaves unchanged the value $x$, regardless of the observed outcome. When proving this result for improper mixtures, Herbut implicitly adopts “composition compatibility” to describe local measurements.
@@
Martinez [undefaai] offers an early account of known derivations of the Lüders rule including Herbut’s work but not Bell and Nauenberg’s contribution. The paper also demonstrates that the converse of Lüders theorem [undefaaf, undefr] implies the quantum mechanical collapse for non-composite systems: if, for any pair of commuting observables $M_{1}$ and $M_{2}$ and any initial state, we assume that the sequence of measurements $M_{1}\prec M_{2}\prec M_{1}$ always produces identical outcomes for $M_{1}$ regardless of the outcome of the $M_{2}$ measurement, then the post-measurement state must necessarily be given by the Lüders rule. Martinez further examines a quantum-logical derivation employing the “Sasaki hook” presented in [undefai]. The extension of the update rule to local measurements on subsystems is not considered.
@@
The proof that establishes the Lüders instrument as the unique ideal quantum instrument for sharp discrete measurements dates back to Davies [undefz], as recognised in [undefr, undefq]. The notion of a quantum instrument incorporates two key assumptions: convex-linearity of the single-system state-update rule, and the standard linear extension to composite systems enabled by the complete positivity of $\omega_{A}$ (see Eq. (18) in Sec. 5.1). This argument for isolating the Lüders collapse for single systems has been rediscovered occasionally (see [undefaah], for example).
@@
In 2019, Masanes et al. advance the argument that both the Born rule and the post-measurement state rule can be derived from the other postulates of quantum theory when supplemented by suitable operational constraints [undefaak]. In doing so, they consider alternative rules for computing outcome probabilities, explicitly discussing both composite systems and local measurements. Having established the uniqueness of the Born rule within their framework, they assert that quantum instruments constitute the only consistent description of state updates. This conclusion, along with their broader framework, has attracted some criticism [undefaaw, undefak, undefax, undefaaj, undefaav]. The argument for quantum instruments depends on an assumption regarding sequential measurements that, as already shown in [undefah], is mathematically equivalent to requiring convex-linearity of state updates.
@@
Conceptually, Masanes et al.’s approach mirrors Kleinmann’s framework [undefaab] in that it is limited a priori to update rules satisfying preparation indistinguishability. This restriction explains why they adopt the completely positive extension (18) for local measurements on composite systems: any other extension would necessarily violate their assumption of preparation indistinguishability.
@@
In retrospect, the Lüders rule is seen to have been derived on the basis of operational, information-theoretic, quantum-logical, and epistemically motivated assumptions. Most of the derivations are limited to single-system update rules, and their extension to local measurements on subsystems is either ignored or tacitly assumed. The extension of state-update rules is, however, not trivial, since the single-system rule alone does not determine it. Additional structural assumptions such as preparation indistinguishability or composition compatibility (cf. Theorem 2) are required to exclude alternative update rules in the multi-partite case.